A certain recipe asks for 3/5 cups of granola.How much is needed to make 1/2 of a recipe?
Found 2 solutions by 428225, fractalier:
Answer by 428225(90)

This question can be solved by creating an equation. You know that for every full recipe you need 3/5 cups of granola; let me create an equation for this: let x equal the amount of granola needed in a full recipe

3/5 = x

3/5 is also equivalent to 6/10

∴ 6/10 = x

Now if you only want half of the recipe, you also have to half your equation

∴ 3/10 = x/2

∴ For half a recipe you need 3/10 cups of granola.

Answer by fractalier(6550)

If the whole recipe calls for 3/5 cups granola, then half a recipe would call for half of 3/5 cups, right?
(1/2)(3/5) = 3/10 cups granola